# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

### v4

@@ 3.1 Convergence of ChFSI with inexact subspace construction
For an n-dimensional space $\mathcal{S}^{(i)}$ satisfying $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ and $\underline{\mathcal{S}}^{(i+1)}=\mathcal{R}\left(\underline{C_{p}(\textbf{H})\textbf{X}^{(i)}}\right)$ we can write
@@
where $\hat{\boldsymbol{\Delta}}_{p}^{(i)}=\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}-C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}$
@@
Partitioning the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$, we have $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl[\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr]$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible. Consequently, we can write
$\hat{\textbf{U}}^{\dagger}\underline{\mathcal{S}}^{(i+1)}$ $=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}\right)=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}+\hat{\textbf{U}}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)$
@@
Consequently, from lemma A.1 we can write
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $=\left\lVert\left(C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)\left(C_{p}(\boldsymbol{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)^{-1}\right\rVert\,$
$\leq\frac{\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}{1-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}$ (4)
Note that in the last step, we have assumed that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,<1$, and using the fact that $\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,=\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}$, a sufficient condition for this to be true is $\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$. Upon further simplification using sub-multiplicative property of matrix spectral norms and triangle inequality, the inequality in eq. 4 can be written as
@@
We note that if $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ remains nearly constant with iteration $i$, the right-hand side of the inequality in eq. 5 keeps increasing as we approach the exact eigenspace and beyond a certain angle $\angle(\mathcal{S}^{(i)},\mathcal{S})$ this inequality gets violated. Consequently, the angle stops decreasing and we cannot approach the exact eigenspace beyond that point. To this end, for robust convergence, we require that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ also decreases as we approach the exact eigenspace, and we demonstrate that our proposed residual-based reformulation of Chebyshev filtered subspace iteration procedure (R-ChFSI) algorithm described subsequently accomplishes this. We further note that $\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\geq 2\sqrt{2}$ and consequently for convergence with approximations, we obtain the following necessary condition on $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ from eq. 5,
@@ 3.2.1 Traditional Chebyshev filtering method
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using inexact matrix-products of the recurrence relation eq. 7 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ and $\tilde{\eta}_{k}$ are some finite constants that depend on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
$\textbf{Z}_{k}^{(i)}=\textbf{B}\textbf{R}_{k}^{(i)}=\textbf{D}(C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)}))=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})\;\;\;\text{for}\;\;\;k=0,\dots,p$ (10)
where we have defined $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ and recall $p$ is the maximum Chebyshev polynomial degree used in the subspace construction step. The motivation behind this definition is the fact that both $C_{k}(\textbf{H})$ and H share the same eigenvectors with the eigenvalues being $C_{k}(\boldsymbol{\Lambda})$ and $\boldsymbol{\Lambda}$ respectively, this ensures that as $\textbf{X}^{(i)}$ and $\boldsymbol{\Lambda}^{(i)}$ approach the exact eigenvector and eigenvalues the weighted residual $\textbf{Z}_{k}^{(i)}$ approaches zero. Now, if we rewrite the recurrence relation defined in eq. 2 such that the matrix products are in terms of $\textbf{Z}_{k}^{(i)}$ then the relative error in the matrix products is then proportional to $\lVert\textbf{Z}_{k}^{(i)}\rVert\,$ thus ensuring a lower absolute error in the filtered subspace. We now propose the following recurrence relation that accomplishes our objective
@@
The recurrence relation given by eq. 2 can be reformulated in terms of the weighted residuals defined by $\textbf{Z}_{k}^{(i)}=\textbf{B}\textbf{R}_{k}^{(i)}=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})$ as
@@
Multiplying with $\textbf{X}^{(i)}$, subtracting from eq. 2, and denoting $\textbf{Z}_{k}^{(i)}=\textbf{B}\textbf{R}_{k}^{(i)}$ we have
@@
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (14)
@@
From lemmas A.4, 3.4, and A.5 we have $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from eq. 3 we require
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (15)
@@
Finally, the Chebyshev filtering step in R-ChFSI can be summarised in the following algorithm.
@@
INPUTS: Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$ and eigenvalues $\boldsymbol{\Lambda}^{(i)}$
OUTPUT: The filtered subspace $\textbf{Y}_{p}^{(i)}$
TEMPORARY VARIABLES:X, Y, $\textbf{R}_{\textbf{X}}$, $\textbf{R}_{\textbf{Y}}$$\boldsymbol{\Lambda}_{\textbf{X}}$ and $\boldsymbol{\Lambda}_{\textbf{Y}}$
@@
We would also like to comment that while our above mathematical analysis is focused on generalized eigenvalue problems, it is trivial to propose a robust mixed precision eigensolver for standard eigenvalue problems by setting $\textbf{D}=\textbf{B}=\textbf{I}$. We note that the mixed precision R-ChFSI method for the standard eigenvalue problem is significantly more robust than the naive mixed-precision ChFSI algorithm, wherein we would perform the matrix products in eq. 2 in lower precision. The naive mixed precision ChFSI algorithm does not converge beyond a certain point, as evident from the discussion following eq. 8 and our numerical studies. In the interest of brevity, the subsequent section of this manuscript will focus solely on solutions for generalised eigenproblems using both low-precision and approximate inverses in computing matrix products for constructing the subspace rich in the desired subspace using Chebyshev filtering.

### v5

@@ 3.1 Convergence of ChFSI with inexact subspace construction
For an n-dimensional space $\mathcal{S}^{(i)}$ satisfying $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ and $\underline{\mathcal{S}}^{(i+1)}=\mathcal{R}\left(\textbf{B}^{\frac{1}{2}}\underline{C_{p}(\textbf{H})\textbf{X}^{(i)}}\right)$ we can write
@@
where $\hat{\boldsymbol{\Delta}}_{p}^{(i)}=\textbf{B}^{\frac{1}{2}}\boldsymbol{\Delta}_{p}=\textbf{B}^{\frac{1}{2}}(\underline{C_{p}(\textbf{H})\textbf{X}^{(i)}}-C_{p}(\textbf{H})\textbf{X}^{(i)})$
@@
Note that $\hat{\textbf{X}}^{(i)}$ is an orthogonal matrix by virtue of implicit B orthogonalization of $\textbf{X}^{(i)}$ in the Rayleigh-Ritz step. Define the $n\times n$ matrix $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and the $(m-n)\times n$ matrix $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. Partitioning the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$, we have $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl[\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr]$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible. Using the maximum principal angle given by definition 2.1, we have $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}\rVert\,$. In the eigenvector coordinate system, the approximate filtered subspace can be written as
@@
Using definition 2.1, we can write
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $=\left\lVert\left(C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)\left(C_{p}(\boldsymbol{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)^{+}\right\rVert\,$
$\leq\frac{\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}{1-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}$ (4)
Note that in the last step, we have assumed that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,<1$, and using the fact that $\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{+}\rVert\,=\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}$, a sufficient condition for this to be true is $\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$. Upon further simplification using sub-multiplicative property of matrix spectral norms and triangle inequality, the inequality in eq. 4 can be written as
@@
We note that if $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ remains nearly constant with iteration $i$, the right-hand side of the inequality in eq. 5 keeps increasing as we approach the exact eigenspace and beyond a certain angle $\angle(\mathcal{S}^{(i)},\mathcal{S})$ this inequality gets violated. Consequently, the angle stops decreasing, and we cannot approach the exact eigenspace beyond that point. To this end, for robust convergence, we require that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ also decreases as we approach the exact eigenspace, and we demonstrate that our proposed residual-based reformulation of Chebyshev filtered subspace iteration procedure (R-ChFSI) algorithm described subsequently accomplishes this. We further note that $\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\geq 2\sqrt{2}$ and consequently for convergence with approximations, we obtain the following necessary condition on $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ from eq. 5,
@@ 3.2.1 Traditional Chebyshev filtering method
The spectral norm of the error $\hat{\boldsymbol{\Delta}}_{k}^{(i)}=\textbf{B}^{\frac{1}{2}}(\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)})$ in the subspace construction using inexact matrix-products of the recurrence relation eq. 7 satisfies $\lVert\hat{\boldsymbol{\Delta}}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ and $\tilde{\eta}_{k}$ are some finite constants that depend on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
$\textbf{Z}_{k}^{(i)}=\textbf{D}\textbf{R}_{k}^{(i)}=\textbf{D}(C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)}))=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})\;\;\;\text{for}\;\;\;k=0,\dots,p$ (10)
where we have defined $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ and recall $p$ is the maximum Chebyshev polynomial degree used in the subspace construction step. The motivation behind this definition is the fact that both $C_{k}(\textbf{H})$ and H share the same eigenvectors with the eigenvalues being $C_{k}(\boldsymbol{\Lambda})$ and $\boldsymbol{\Lambda}$ respectively, this ensures that as $\textbf{X}^{(i)}$ and $\boldsymbol{\Lambda}^{(i)}$ approach the exact eigenvectors and eigenvalues the weighted residual $\textbf{Z}_{k}^{(i)}$ approaches zero. Now, if we rewrite the recurrence relation defined in eq. 2 such that the matrix products are in terms of $\textbf{Z}_{k}^{(i)}$ then the relative error in the matrix products is then proportional to $\lVert\textbf{Z}_{k}^{(i)}\rVert\,$ thus ensuring a lower absolute error in the filtered subspace. We now propose the following recurrence relation that accomplishes our objective
@@
The recurrence relation given by eq. 2 can be reformulated in terms of the weighted residuals defined by $\textbf{Z}_{k}^{(i)}=\textbf{D}\textbf{R}_{k}^{(i)}=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})$ as
@@
Multiplying with $\textbf{X}^{(i)}$, subtracting from eq. 2, and denoting $\textbf{Z}_{k}^{(i)}=\textbf{D}\textbf{R}_{k}^{(i)}$ we have
@@
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,(\gamma_{m}\eta_{p}+\zeta\tilde{\eta}_{p})\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (14)
@@
From lemmas A.3, 3.4, and A.4 we have $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,(\gamma_{m}\eta_{p}+\zeta\tilde{\eta}_{p})\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from eq. 3 we require
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,(\gamma_{m}\eta_{p}+\zeta\tilde{\eta}_{p})\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (15)
@@
Finally, the Chebyshev filtering step in R-ChFSI can be summarized in the following algorithm.
@@
INPUTS: Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$, the initial guess of eigenvectors $\textbf{X}^{(i)}\in\mathbb{C}^{m\times n}$, current eigenvalue estimates $\boldsymbol{\Lambda}^{(i)}=\mathrm{diag}(\epsilon_{1}^{(i)},\dots,\epsilon_{n}^{(i)})$, approximate inverse of B denoted $\textbf{D}^{-1}$, and the residual $\textbf{Y}=\textbf{A}\textbf{X}^{(i)}-\textbf{B}\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$.
OUTPUT: The filtered subspace $\textbf{Y}_{p}^{(i)}\in\mathbb{C}^{m\times n}$
TEMPORARY VARIABLES: $\textbf{R}_{\textbf{X}}\in\mathbb{C}^{m\times n}$, $\textbf{R}_{\textbf{Y}}\in\mathbb{C}^{m\times n}$ (filtered residual vectors), $\boldsymbol{\Lambda}_{\textbf{X}}\in\mathbb{R}^{n\times n}$, $\boldsymbol{\Lambda}_{\textbf{Y}}\in\mathbb{R}^{n\times n}$ (Chebyshev-filtered eigenvalue matrices)
@@
We would also like to comment that while our above mathematical analysis is focused on generalized eigenvalue problems, it is trivial to extend R-ChFSI to standard eigenvalue problems by setting $\textbf{D}=\textbf{B}=\textbf{I}$. When $\textbf{D}^{-1}=\textbf{B}^{-1}$ (i.e. the approximate inverse is exact) and the same matrix is used for both the Chebyshev filter and the Rayleigh–Ritz projection, ChFSI and R-ChFSI are algebraically equivalent. However, when the matrix–vector product in the filter is inexact—e.g. the filter applies an approximate $\tilde{\textbf{A}}$ while the Rayleigh–Ritz step uses the exact A—R-ChFSI retains its robustness advantage even in the standard eigenproblem. The same advantage arises from a mismatch $\textbf{D}^{-1}\neq\textbf{B}^{-1}$ in the generalized eigenproblem, which is precisely the scenario encountered in practice where forming an exact $\textbf{B}^{-1}$ is prohibitively expensive. In the interest of brevity, the subsequent section of this manuscript will focus on controlled experiments validating these claims on dense random matrices of moderate size, followed by solutions for generalized eigenproblems from DFT using both low-precision and approximate inverses in computing matrix products for constructing the subspace rich in the desired subspace using Chebyshev filtering.